A bearing is selected a random and measured. The radius of ball bearings


produced by a particular process follow a normal distribution with y =


0.255 and g=0.001 inches. What is the probability that the radius is at


most 0.2545 in?


1
Expert's answer
2022-05-17T15:03:08-0400
"P(X\le0.2545)=P(Z\le\dfrac{0.2545-0.255}{0.001})"

"=P(Z\le -0.5)\approx0.3085"